# Generative Omnimatte: Learning to Decompose Video into Layers

### v1

@@ 2 Related Work
Video Generation and Inpainting. Recent advancements in diffusion models have significantly improved the capabilities of text-to-video generation. These models [6, 7, 11, 58, 19, 46, 15, 56, 63, 54] are trained on massive datasets of video-text pairs, enabling them to generate visually appealing videos from text descriptions.
@@
Object Effect Removal. Work [30, 12] in video shadow removal focuses specifically on removing cast shadows, but does not aim to perform object-shadow association, object removal, or background completion. Recently, ObjectDrop [51] finetunes an image inpainting model for object and effect removal. However, it is unsuitable for video object removal due to a lack of temporal coherence when run on individual frames. Moreover, our goal extends beyond object removal; we aim to decompose a video into an entire set of omnimattes to enable layer-based video editing tasks.
@@ 5 Discussion and Limitations
We provide additional details on the training process of our object-effect-removal model (Sec. A), the inference pipeline for object and effect removal (Sec. B), the optimization details for omnimatte reconstruction (Sec. C), the quantitative comparisons (Sec. D), and the quantitative ablation study (Sec. E). The video comparisons, training video examples, and self-attention visualization are provided in the project page.
Appendix A Training Details of Our Casper Model
@@
We finetune our Casper model from the pretrained Lumiere inpainting model [6, 36] for 20,000 iterations and a batch size of 32.
Appendix B Inference Details of Object and Effect Removal
@@
Appendix C Omnimatte Optimization Details
@@
Appendix D Onimatte-RF Synthetic Evaluation Benchmark
@@
Appendix E Ablation study

### v2

@@ 2 Related Work
Video Generation and Inpainting. Recent advancements in diffusion models have significantly improved the capabilities of text-to-video generation. These models [7, 8, 16, 12, 60, 21, 48, 17, 58, 65, 56] are trained on massive datasets of video-text pairs, enabling them to generate visually appealing videos from text descriptions.
@@
Object Effect Removal. Work [32, 13] in video shadow removal focuses specifically on removing cast shadows, but does not aim to perform object-shadow association, object removal, or background completion. Recent methods [53, 4] fine-tune image diffusion models for object and effect removal. However, they are unsuitable for video object removal because they lack temporal coherence when run on individual frames. Moreover, our goal extends beyond object removal; we aim to decompose a video into an entire set of omnimattes to enable layer-based video editing tasks.
@@ 5 Discussion and Limitations
Acknowledgments. We would like to thank Charles Herrmann, Daniel Geng, Junhwa Hur, Hadi Alzayer, Deqing Sun, Daniel Vlasic, Miki Rubinstein, and Bill Freeman for their insightful discussions and generous support. We sincerely appreciate Ray Li, Scott Wisdom, Andrei Kapishnikov, Matthew Grimes, AJ Maschinot, and Shu-Jung Han for their invaluable help in filming testing videos and contributing creative editing ideas.
@@
We provide additional discussions (Sec. A), as well as further details on the training process of our object-effect-removal model (Sec. B), the inference pipeline for object and effect removal (Sec. C), the optimization details for omnimatte reconstruction (Sec. D), runtime analysis (Sec. E), the quantitative comparisons (Sec. F), the quantitative ablation study (Sec. G). The video comparisons, training video examples, and self-attention visualization are provided in the project page.
Appendix A Additional Discussions
Reproducibility.
We will release our dataset for reproducibility. In addition, we finetuned a publicly available CogVideoX [56] using the same data to create a DiT version of Casper [^2] For CogVideoX-based Casper finetuning, we adopt the codes and model from the third-party github: https://github.com/aigc-apps/CogVideoX-Fun.git, without adjusting the original CogVideoX hyperparameters. The results are shown in Fig. 13. The sampling process takes 66 sec for an 85-frame, 384$\times$672 video and a trimask with 50 DDIM steps w/o CFG on an A100 GPU.
Handling occluded foreground.
Impact of text prompts.
We used simple, short prompts such as “a clean beach” for fine-tuning the Lumiere-based Casper. While Casper is primarily driven by the input video and trimask and less sensitive to text prompts (“empty scene” works well for both rows below), we did observe that a highly unrelated prompt (e.g., “rock concert”) could affect the performance, as shown in Fig. 15.
Figure 13: Object-effect-removal results of CogVideoX-based Casper.
Figure 15: Impact of the text prompt.
Challenges in handling similar objects.
While our Kubric dataset includes multiple similar objects to help Casper handle such cases, it may still struggle with complex scenarios. We observed a similar issue in CogVideoX-based Casper, potentially due to the domain discrepancies between Kubric data and real-world videos. This could be mitigated in future work through more realistic data generation and re-introducing successful removal results into the training set.
Undesired detail changes.
These artifacts are caused by Lumiere’s SSR model, which can hallucinate high-frequency details. While we use a post-processing step to transfer original details, it is applied conservatively to avoid altering effect-removal areas. A future direction may be exploring an SSR model that can refer to the original video when upsampling the removal result.
Appendix B Training Details of Our Casper Model
@@
Object-Paste data.
Due to the challenge of real data collection, Object-Paste is synthesized to enhance inpainting and background preservation, but not effect removal. Though not always reflected in metrics, we observed qualitative improvements in the main paper Fig. 9a and Supp. HTML. A tradeoff exists between background preservation and effect removal, and a high Object-Paste weight may hinder effect-removal performance. To balance this, we set its weight to 2% in the training set.
@@
We fully finetune our Casper model from the pretrained Lumiere inpainting model [7, 38] for 20,000 iterations and a batch size of 32.
Appendix C Inference Details of Object and Effect Removal
@@
Appendix D Omnimatte Optimization Details
@@
Appendix E Runtime
For Stage 1, Casper takes 12 min to process an 80-frame video and 15-20 min for longer videos, the same as the original Lumiere-Inpainting. We run Casper on a 96GB TPU with a batch size of 4 (e.g., three solo videos and a clean background). Additional objects can be run in parallel with multiple TPUs. For Lumiere SSR upsampling, it takes around 15 minutes. For Stage 2, our unoptimized code takes 7 min to produce each object layer on a 48GB TPU. Each object layer is computed independently and thus can be parallelized. Once obtaining all layers, the post-processing detail transfer takes 1 minute. For an 80-frame video of 3 objects on a single TPU, the entire process takes $12+15+7\times 3+1$ = 49 min (or $12+15+7+1=35$ min with multi TPUs). In contrast, OmnimatteRF takes 3 hr to optimize a NeRF and all layers together, potentially limiting the number of objects due to GPU memory constraints.
Appendix F Onimatte-RF Synthetic Evaluation Benchmark
@@
Appendix G Ablation study